# Audio-Infused Automatic Image Colorization by Exploiting Audio Scene Semantics

### v1

@@ III-E Colorization backbone
| InstColor [10] | 0.1423 | 25.4754 | 0.9529 | 0.1365 | 25.2331 | 0.9601 |
| ChromaGAN [30] | 0.1423 | 24.5931 | 0.9414 | 0.1337 | 25.0703 | 0.9554 |
| Deoldify [52] | 0.1566 | 24.2527 | 0.9375 | 0.1465 | 24.8899 | 0.9563 |
| BigColor [12] | 0.1587 | 22.9913 | 0.8957 | 0.1542 | 22.9276 | 0.8988 |
| DisColor [33] | 0.1508 | 24.3153 | 0.9487 | 0.1367 | 24.6226 | 0.9592 |
| Ours | 0.1282 | 25.6939 | 0.9521 | 0.1231 | 25.4528 | 0.9610 |
IV Experiments
IV-A Dataset
Considering that there is no audio in the existing colorization datasets, we collect an audiovisual colorization (AVColor) dataset, mainly from VGGSound [53], which contains a lot of 10s videos with synchronized audio and images. Unlike other audiovisual multimodal tasks, such as localization and recognition, the task in this study has some specificity in that the audio should convey some visual scene color information, e.g., an underwater bubbling sound corresponding to the water-color scene and a ski sound corresponding to the snow-color scene. Therefore, in order to investigate the potential assistance of audio in colorization, we established the AVColor dataset consisting of 13 categories for training and testing. These categories are underwater bubbling, rooster crowing, raining, skiing, footsteps on snow, fire crackling, thunder, chainsawing trees, lawn mowing, fire truck siren, chopping wood, chicken clucking and sea waves.
For each video, we select one frame and form an audio-image pair with the corresponding audio. The final training set (AVColor_Train) contains 5383 pairs and the test set (AVColor_Test1) contains 130 pairs. In addition, to verify the generalization ability of our method, from other existing audiovisual datasets [54, 17, 55], we construct another test set, denoted as AVColor_Test2, containing 50 audiovisual pairs, where the rooster image is selected from animal10 [56] and the corresponding sound is selected from ESC-50 [57].
IV-B Implementation details
For the colorization backbone, we initialize the and modules with the pretrained weights provided by [33]. For the visual scene semantic extraction module , we use the existing classification network VGG-19 [48], which is selected before the penultimate layer. The pretrained Resnet-18 on VGGSound is used as the audio feature extraction network . During the training process, we freeze the parameters of and , and set , . We set the batch size to 16 and minimize the objective loss using Adam optimizer with , . All the images are rescaled to 256256. Source codes and the dataset will be released.
IV-C Comparison with existing methods
| Settings | AVColor_Test1 | AVColor_Test2 | ||||
| LIPIS↓ | PSNR↑ | SSIM↑ | LIPIS↓ | PSNR↑ | SSIM↑ | |
| full model | 0.1282 | 25.6939 | 0.9521 | 0.1231 | 25.4528 | 0.9610 |
| w/o fine-tuning | 0.1475 | 25.8181 | 0.9483 | 0.1546 | 25.564 | 0.9580 |
| w/o learing | 0.1513 | 24.4861 | 0.9409 | 0.1466 | 23.725 | 0.9462 |
| w/o vc_pretrain | 0.1386 | 25.1216 | 0.9472 | 0.1430 | 24.163 | 0.9523 |
| w/o audio ( baseline ) | 0.1508 | 24.3153 | 0.9487 | 0.1367 | 24.6226 | 0.9592 |
Quantitative comparisons. We compare our proposed method with recent automatic colorization methods, including InstColor [10], ChromaGAN [30], Deoldify [52], BigColor [12], and DisColor [33], where for ChromaGAN, we directly employ labels for coloring assistance in the training phase, and BigColor takes the category information as the input. By contrast, our method does not require labels during either the training or testing stages. As shown in Table I, our method performs better in terms of these metrics, indicating that the quality of our generated image is better and more similar to the color of the original image. At the same time, our proposed method has certain advantages when tested directly on the second test set, indicating that the model has some generalization ability.
Qualitative comparisons. Qualitative comparisons are performed to demonstrate the effectiveness of the proposed method. From Fig. 5 we find that by relying only on a single visual modality, the existing visual models are sometimes unable to find the correct semantic information of the scene, making the generated color not match the actual situation. For example, for the snow image in the third row, existing methods tend to yield a greenish or purple color owing to fewer contextual clues. In fact, this grayscale image corresponds to the scene of skiing. When we inject the counterpart ski sound, we can find that it can complement the scene knowledge for the model and correct the generated color. The same is true for the diving scene shown in the first row. For images in which the overall color is not distinct, the associated sound could still enhance the color of the subjects in the scene, such as the color depth of the red crown of a rooster in the second row.
IV-D Ablation study
The role of audio. First, we compare with the grayscale image colorization model based on a single visual modality, i.e., the setting ‘w/o audio (baseline)’ in Table II. As can be seen in Fig. 6, the addition of relevant audio leads to a better understanding of scenes. For example, in the scene of walking on snow, from the grayscale image alone, the baseline model would colorize the leaves with common green color and thus lead to wrong colorization. Besides, the red color of the fire truck body is deepened by the addition of the fire siren sound. The scores in Table II also show that audio can make the generated results more consistent with the actual scene.
@@ IV-D Ablation study
Audio scene semantic learning. If we incorporate audio directly into the visual model for end-to-end training, i.e., ‘w/o leaning’, due to the modal heterogeneity between audio and vision, the model usually ignores the role of audio and cannot successfully establish the correspondence between audio and colorization. From Table II, we find that the direct addition of audio even decreases the metrics compared to the baseline. In addition, Fig. 6 shows that the colors of the waves and the water are completely incorrect, which demonstrates the importance of scene semantic learning of audio.
Guidance of real scene semantic priors. When the learned audio potential scene semantic features are directly fed into the baseline for end-to-end training, i.e., the setting ’w/o vc_pretrain’ in Table II, we find that although the color semantics generated in the grayscale image regions related to the sound is relatively accurate, such as the rooster crow in Fig. 7, the overall visual subjective effect becomes worse in this case. For example, the green color of the grass is very unnatural and there exists color bleeding. When pre-training with real scene semantic priors, the subjective effect of color is greatly improved. It shows that it is easier to learn the correspondence between semantics and color reasoning in the same visual modality.
Fine-tuning. Here we directly replace the real semantic feature with the audio scene feature for inference, i.e., ’w/o fine-tuning’. From the colors of the fire truck and the crown of the rooster in Fig. 7, it can be seen that does contain the scene color information, but the overall coloring effect is not very satisfactory at this time, and the same conclusion is obtained from the comparison of the metrics in Table II. After further fine-tuning, it makes the colors of the image scenes more realistic and pleasing.
IV-E Analysis of audio scene semantics
We analyze the audio scene semantics to determine the color information to which the sound actually responds in the scene. We keep the input image unchanged and swap the input audio. Then, by observing the color after exchanging the audio, we can laterally obtain the scene color corresponding to the current input audio. From Fig. 8, we can find that the sound of the rooster corresponds to the red semantic of the crown; the sound of a ski corresponds to the white of the snow; the sound of bubbling underwater corresponds to the blue of the deep sea; the siren corresponds to the red of the ambulance; the sound of the lawn mower corresponds to the green of the grass, and the sound of footsteps in the snow corresponds to white.
These results demonstrate the ability of our network to learn the scene color semantics of the input audio, thereby enabling the decoupling of audio-visual multimodal features and enhancing the understanding of the scenes.
IV-F Failures and limitations
As shown in Fig. 9, the proposed method sometimes causes scene colors to overflow or suppress the colors of other objects. In future work, we tend to build larger training set with high-resolution images to further improve the performance of the proposed method.
Furthermore, our method is predominantly suitable for scenarios characterized by a strong correspondence between audio and video signals. When the audio and visual content lack correlation, such as in instances involving voiceovers or background music found in old movies, significant color deviations can arise, especially in some grayscale scenes where the visual semantic is not obvious (e.g., snow field). In such cases, how to maintain the backbone performance of image colorization deserves future research.
V Conclusion
@@
This paper proposes an audio-guided automatic image colorization method for the first time, which can use corresponding audio information to enhance the scene semantics and improve the colorization performance. The network is trained in three steps without manual labels of audio semantics. First, the colorization backbone is pretrained with scene semantics extracted from the visual domain. Then, the optimized visual scene semantics are adopted to constrain the learning of audio semantics. Finally, the audio semantics are used to improve the coloring process. Moreover, we build an audiovisual colorization dataset for training and testing. Experimental results demonstrate that the proposed audio-guided method can outperform other state-of-the-art methods.
Acknowledgments
@@
This work is supported by the National Natural Science Foundation of China under Grant 61972127, Grant 61972129, and Grant 62076086. And it is partly supported by the grants of the Key Research and Development Program in Anhui Province 202004d07020008.

### v2

@@ 4.2 Results of Colorization
full (DisColor-based) | 0.147 | 24.463 | 0.938 | 0.136 | 25.475 | 0.946 |
w/o multi-step | 0.163 | 23.752 | 0.929 | 0.151 | 24.583 | 0.932 |
[…]
w/o (missing audio) | 0.807 | 7.467 | 0.103 | 0.792 | 7.670 | 0.098 |
[…]
full (DDC-based) | 0.138 | 24.446 | 0.936 | 0.128 | 25.315 | 0.941 |
w/o multi-step | 0.144 | 24.200 | 0.930 | 0.131 | 25.128 | 0.938 |
w/o | 0.143 | 24.174 | 0.929 | 0.131 | 25.089 | 0.933 |
[…]
w/o multi-step | 0.153 | 24.471 | 0.922 | 0.137 | 25.235 | 0.927 |
[…]
w/o (missing audio) | 0.639 | 10.546 | 0.311 | 0.628 | 10.757 | 0.292 |
full (missing audio) | 0.157 | 24.490 | 0.920 | 0.142 | 25.319 | 0.923 |
4.3 Ablation Study
Effectiveness of audio. To further explore the effectiveness of audio, we directly exclude audio by setting r to 0 in Eq. 8. As illustrated in Table 2, comparing ‘full’ and ‘full (missing audio)’, we observe a noticeable performance decline when audio is omitted, which shows that audio can effectively improve colorization performance. Additionally, Fig. 5 (b) provides some qualitative comparisons. It can be found that the addition of relevant audio leads to a better understanding of scene; for instance, blue for diving, red for ambulance.
@@ 4.3 Ablation Study
Effectiveness of multi-step training. The purpose of the multi-step training is to learn the implicit scene color semantics of the audio, thus providing an effective aid for visual coloring. If we incorporate audio directly into the visual model for end-to-end training, i.e., ‘w/o multi-step’, due to the modal heterogeneity between audio and vision, the model usually ignores the role of audio and cannot successfully establish the correspondence between audio and colorization. Fig. 5 (a) shows that the colors of the sky and snow are completely incorrect, which demonstrates the importance of scene semantic learning of audio.
Effectiveness of Relevance Mechanism (RM) in the DSG module. The RM is designed to enhance the robustness of the AIAIC model. Specifically, it can increase dependency on the visual backbone for colorization when audio and vision are irrelevant. Moreover, when audio is not available, degradation to a visual-only model can still allow for basic colorization. To validate these, we conduct two corresponding ablation experiments. 1) Comparison between ‘full’ and ‘w/o ’ settings in Table 2 demonstrates that incorporating the RM generally improves performance, especially on AVE dataset containing irrelevant segments. 2) Furthermore, when audio is unavailable, i.e., ‘w/o (missing audio)’, colorization fails entirely, as indicated by significant discrepancies in the LIPIS and SSIM metrics compared to the ‘full’ setting. Conversely, in the setting ‘full (missing audio)’, adding this mechanism allows the model to sustain a certain level of colorization effect as in the visual baseline.
5 Conclusion
@@
This paper proposes a novel and pluggable audio-infused automatic image colorization method for the first time, which can use corresponding audio information to enhance the scene semantics and improve the colorization performance. The network is trained in three steps without manual labels of audio semantics. First, the colorization backbone is pretrained with scene semantics extracted from the visual domain. Then, the optimized visual scene semantics are adopted to constrain the learning of audio semantics. Finally, the audio semantics are used to improve the coloring process. Experimental results demonstrate the effectiveness of our proposed audio-guided method.
5.0.1 Acknowledgements
@@
This work is supported by the National Natural Science Foundation of China under Grant 61972127, Grant 61972129.